1.      Given that csc A = , A in QI, and sec B = , sin B < 0, find

a.      cos (A – B)

b.     tan (A – B)


1
Expert's answer
2021-03-04T06:02:57-0500
"\csc A=\dfrac{1}{\sin A}=a, a\geq 1"


"\sin A=\dfrac{1}{a},"

"\cos A=\sqrt{1-\sin^2 A}=\sqrt{1-(\dfrac{1}{a})^2}=\dfrac{\sqrt{a^2-1}}{a}"

"\sec B=\dfrac{1}{\cos B}=b, |b|\geq 1"

"\cos B=\dfrac{1}{b}"

"\sin B=-\sqrt{1-\cos^2 B}=-\sqrt{1-(\dfrac{1}{b})^2}=-\dfrac{\sqrt{b^2-1}}{|b|}"

a.


"\cos(A-B)=\cos A\cos B+\sin A\sin B"

"=\dfrac{\sqrt{a^2-1}}{a}(\dfrac{1}{b})+\dfrac{1}{a}(-\dfrac{\sqrt{b^2-1}}{|b|})"

"b\geq1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}-\sqrt{b^2-1}}{ab}"


"b\leq-1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}+\sqrt{b^2-1}}{ab}"

b.


"\tan(A-B)=\dfrac{\sin(A-B)}{\cos(a-B)}"

"\sin(A-B)=\sin A\cos B-\cos A\sin B"

"=\dfrac{1}{a}(\dfrac{1}{b})-\dfrac{\sqrt{a^2-1}}{a}(-\dfrac{\sqrt{b^2-1}}{|b|})"

"b\geq1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}-\sqrt{b^2-1}}{ab}"


"\sin(A-B)=\dfrac{1+\sqrt{(a^2-1)(b^2-1)}}{ab}"

"\tan(A-B)=\dfrac{1+\sqrt{(a^2-1)(b^2-1)}}{\sqrt{a^2-1}-\sqrt{b^2-1}}"


"b\leq-1, a\geq1:\cos(A-B)=\dfrac{\sqrt{a^2-1}+\sqrt{b^2-1}}{ab}"


"\sin(A-B)=\dfrac{1-\sqrt{(a^2-1)(b^2-1)}}{ab}"

"\tan(A-B)=\dfrac{1-\sqrt{(a^2-1)(b^2-1)}}{\sqrt{a^2-1}+\sqrt{b^2-1}}"